Do not assume a constant g for this problem. The gravitational potential energy of a 1 kg object on the earth's surface is -6.28e7 J. How much work is required to lift that object to a height of 5e6 m? GM = 4e14m^3/s^2. Radius of earth = 6.4e6 m.
Answer #1

Initial gravitational potential energy,
U1 = -6.28*10^7 J

Final gravitational potential energy,
U2 = -G*Me*m/(Re + h)

= -6.67*10^-11*5.98*10^24*1/(6.4*10^6 + 5*10^6)

= -3.50*10^7 J

Workd required, W = U2 - U1

= -3.5*10^7 - (-6.28*10^7)

= 2.78*10^7 J <<<<<<<<<<<------------------Answer